# Asymptotics of spherical dynamos exhibiting a small-scale MAC balance

### v1

@@ 4.1 Parameter regime and dependence on the magnetic Prandtl number
As the force balance is known to vary with $Pm$, we start by showing the radial fluctuating force balance as a function of $Pm$ for the case $Ek=10^{-5}$ and $Ra=1.5\times 10^{8}$ in figure 1(a). The force balance for a non-magnetic case is shown to the left of the dashed line. At small values of $Pm$, the force balance for the dynamo cases is similar to the non-magnetic case. As expected, the Lorentz force increases with $Pm$ (Dormy, 2016; Schwaiger et al., 2019; Menu et al., 2020). At the largest value of $Pm=5$ reached here, the Lorentz force is slightly larger than the Coriolis force. The buoyancy force also increases with $Pm$, although less strongly than the Lorentz force – the reason for this increase is discussed later. The remaining terms tend to decrease as $Pm$ is increased, which is likely due to a decrease in flow speeds as the dynamo becomes more efficient at converting kinetic energy to magnetic energy. It is therefore apparent that the dynamo cases have force balances ranging from QG to MAC depending on the value of $Pm$, which suggests that $Pm$ can affect the asymptotic order that the various forces enter in the momentum equation. In this paper, we will fix $Pm=2$, which leads to cases with a MAC force balance, especially at small Ekman numbers. Thus, these results can be contrasted with prior work on the asymptotics of small magnetic Prandtl number dynamos, which are essentially QG (Yan & Calkins, 2022a).
@@ 4.2 Flow speeds
As indicated in figure 2(b), the dynamo cases are described well by the $O(Ek^{-1/3})$ scaling predicted from QG theory. The non-magnetic data appears to be less well described by the $O(Ek^{-1/3})$ scaling. We also plot the rescaled Reynolds number for the non-magnetic cases as a function of $Ra/Ra_{c}$ in figure 2(c), where a better collapse is found. Here, values of $Ra_{c}$ are taken from Christensen & Aubert (2006). The better collapse in this view of parameter space may be due to the slow rate of convergence of $Ra/Ra_{c}$ to the predicted $Ek^{-4/3}$ scaling, as has been noted in linear studies of spherical convection (Jones et al., 2000; Barik et al., 2023). Noting the near linear trend of the convective Reynolds number for the dynamo cases in figure 2(b), we plot the convective Reynolds number compensated by $\widetilde{Ra}\vphantom{Ra}^{-1}$ in figure 2(d). The compensated data has much less dependence on $\widetilde{Ra}$, at least for sufficiently large values of $\widetilde{Ra}$. This scaling can be derived by balancing the Coriolis and buoyancy forces such that
@@
The reduced zonal Reynolds number is shown in figure 2(e), where again the dynamo data appears better collapsed in comparison to the non-magnetic data. If we again assume a balance between the Coriolis and buoyancy forces, but now in the mean momentum equation, then we obtain an identical scaling to that given above for the convective Reynolds number, i.e. $\overline{Re}_{\phi}\sim EkRa/Pr$. The compensated data given in figure 2(f) indicates that the dynamo cases are well described with this scaling, whereas the non-magnetic cases are not. The mean force balance in dynamos is known to be well described by the thermal wind balance (Aubert, 2005; Calkins et al., 2021). However, for non-magnetic convection it is possible for the mean flows to be in geostrophic balance for sufficiently small Ekman number so that only large-scale viscous diffusion is available to saturate the zonal flow, as shown for non-magnetic convection using stress-free mechanical boundary conditions (Nicoski et al., 2024).
@@ 4.3 Force scalings
The rescaled Lorentz force is shown in figure 6(a), where an empirical scaling of $O(Ek^{-3/2})$ has been used to collapse the data. The ratio of the Lorentz force to the Coriolis force is shown in figure 6(b), where it can be seen that this ratio increases with decreasing $Ek$. However, note that while the Lorentz force is scaling more strongly with Ekman number than the Coriolis force, the Coriolis force is larger than the Lorentz force for most of the cases. This observation suggests that as the Ekman number becomes smaller there is either a change in scaling, or that the Lorentz force would become larger than the Coriolis force. It is important to recall that $Pm$ is constant in these cases; as shown in figure 1, the strength of the Lorentz force depends strongly on $Pm$. We also note that this scaling of the Lorentz force together with the scaling of the length scale of the Lorentz force of $O(Ek^{1/6})$ (shown later) implies that the magnetic field scales as $O(Ek^{-1/6})$, which is different than the order one scaling proposed in Calkins et al. (2021). A plot of the rms of the fluctuating magnetic field is provided in figure 6(c), and the corresponding rescaled data is given in figure 6(d).
@@
The temporal behaviour of the fluctuating forces is shown in figure 7 for both a non-magnetic and dynamo case at $Ek=10^{-5}$, $Ra=1.5\times 10^{8}$ ($\widetilde{Ra}=32.3$) at a fixed point in the equatorial plane that lies midway between the shells. In our non-dimensionalisation, time is in units of the large-scale viscous diffusion time, $t_{\nu}$. However, we choose to plot the data in terms of the small-scale viscous diffusion time $Ek^{-2/3}t_{\nu}$, which is the dynamically relevant timescale for QG convection (e.g. Oliver et al., 2023, 2025). For the non-magnetic case, the leading-order force balance is between the Coriolis force and the pressure gradient force, as would be expected for QG convection. The dynamo case is well described by a MAC balance in which the Coriolis force, pressure force, buoyancy force, and Lorentz force are all comparable and leading-order in the dynamics. Although these four terms tend to be the largest in a time averaged sense, the force balance exhibits significant changes with time. For example, at time $Ek^{-2/3}t_{\nu}\approx 6$, there is an approximate balance between the Lorentz and pressure forces, while at other times the Coriolis force is larger than the Lorentz force. Therefore, the leading-order force balance for the non-magnetic cases and the dynamo cases is different, as suggested by the asymptotic scalings for these forces. We note that other studies find the force balance might depend on length scale (e.g. Schwaiger et al., 2019; Aubert et al., 2017; Tassin et al., 2021), with many studies finding that the QG balance holds only at large length scales for dynamo simulations. However, the data shown here indicates that the pointwise balance is not geostrophic.
Further information on the spatial variation of these forces is provided in figure 8, which shows the radial fluctuating forces as a function of latitude for the same cases as in figure 7. For the non-magnetic case, a dominant balance between the Coriolis force and pressure gradient force exists across almost all latitudes. The advection term is larger than the viscous term outside the tangent cylinder, whereas the opposite is true inside the tangent cylinder. This difference is likely due to the higher Rayleigh numbers needed for convection within the tangent cylinder (e.g. Gastine & Aurnou, 2023). For the dynamo case, a dominant balance between the Coriolis force, pressure gradient force, Lorentz force, and buoyancy force prevails across all latitudes, except near the poles where the Coriolis force becomes small. Therefore, for both the dynamo and non-magnetic cases, the dominant force balance does not vary strongly with latitude, although the absolute sizes of the various terms are sensitive to latitude. Most notably, there is a decrease in the size of the Coriolis, pressure gradient, viscous, and buoyancy forces by a few orders of magnitude inside the tangent cylinder for the non-magnetic case.
@@
Defining $\tau$ as the relevant dynamical timescale and balancing inertia with either of these terms then gives
@@
This finding implies that the dynamics of MAC-regime dynamos can be loosely characterised as nonlinear convective Rossby waves.
@@ 4.4 Temperature and dissipation relations
Analysis of the forces in section 4.3 shows that the fluctuating buoyancy force, and therefore the fluctuating temperature, is characterised by a unique asymptotic scaling depending on whether the magnetic field is present. Figure 9 shows the rms of the fluctuating temperature for all cases, where this difference can be observed – for non-magnetic rotating convection the magnitude of the fluctuating temperature scales according to $Ek^{1/3}$, whereas the dynamo cases exhibit a fluctuating temperature that becomes independent of both $Ek$ and $\widetilde{Ra}$, as has been noted previously (Calkins et al., 2021). This difference can be explained based on energetic arguments as follows. For the dynamo cases, a dissipation equation can be derived by dotting the full momentum equation with the velocity field and averaging over both the fluid volume ($V$) and time to give
$\int_{V}$ $u_{r}f_{b}\,dV=\int_{V}u^{\prime}_{r}f^{\prime}_{b}+\overline{u}_{r}\overline{f}_{b}\,dV=\int_{V}\boldsymbol{\omega}^{2}\,dV$ (22)
@@
$\langle u^{\prime}_{r}f^{\prime}_{b}\rangle\sim\langle\boldsymbol{\omega}^{2}\rangle=\langle\mathbf{u}\cdot\mathbf{f}_{v}\rangle,$ (23)
@@ 4.5 Length scales
In this section we compare various length scales for both the dynamo and non-magnetic cases. Figure 10 shows visualisations of both the velocity field and the fluctuating temperature field. As noted in previous work (e.g. Yadav et al., 2016; Schwaiger et al., 2019), the dynamo cases can form length scales that are larger than those observed in comparison to the non-magnetic cases, although small length scales are still present in the dynamo cases. For example, localised small-scale fluid structures with significant flow speeds can be seen in figure 10(b), especially near the inner and outer boundaries. This observation suggests that the dynamo simulations have a wider range of energetically relevant length scales than the non-magnetic cases. Note also the general lack of axial alignment in the dynamo cases, which likely results from the Lorentz force breaking the balance between the Coriolis force and the pressure gradient force.
@@
The linear asymptotic theory of rapidly rotating convection shows that, in addition to the $Ek^{1/3}$ ‘fast’ scale, a ‘slow’ radial scale is also present with an asymptotic scaling of $Ek^{2/9}$ and $Ek^{1/6}$ for the spherical shell and full sphere geometries, respectively (Jones et al., 2000; Dormy et al., 2004). Thus, it might be anticipated that different regions of the spectra should be characterised by different asymptotic scalings. For our investigated range of parameters there is very little difference in these two scaling exponents so that $Ek^{1/6}$ was used here for simplicity. As shown later, this larger length scale may be pertinent to the generation of the magnetic field.
@@
A comparison of the velocity Taylor microscale for the dynamo and non-magnetic cases is shown in figure 13(c). The length scale for the non-magnetic cases tends to be slightly smaller than for the dynamo cases, as might be expected based on the visualisations in figure 10. However, the difference in the Taylor microscale between the non-magnetic and dynamo cases is small, which might be due to viscous dissipation in the dynamo simulations being concentrated in small, localized regions. In addition, the Taylor microscale scales approximately as $O(Ek^{1/3})$ as shown in figure 13(d), although there is still a weak trend in the rescaled data for both the dynamo and non-magnetic cases. This trend might suggest a slow rate of convergence to the predicted QG scaling. Other studies have also found that the velocity field retains an $O(Ek^{1/3})$ length scale (King & Buffett, 2013), though we stress here that this scaling does not necessarily indicate that the viscous force is dominant since, as we have shown, the same length scale can be deduced from the advection term.
@@
Finally, two magnetic length scales are shown in figure 14, where both the magnetic Taylor microscale ($\lambda^{\prime}_{b}$) and a length scale based on the Lorentz force ($\ell^{\prime}_{b}$) are shown. These length scales are defined in equations (11) and (13). Both length scales are comparable in magnitude over the investigated range of parameters and both follow similar trends with $\widetilde{Ra}$ and $Ek$. In particular, for smaller values of $\widetilde{Ra}$ these length scales show an increase up to $\widetilde{Ra}\sim 20-30$, and then a monotonic decrease is observed as $\widetilde{Ra}$ is increased. Figure 14(b) shows the magnetic length scales rescaled with $O(Ek^{1/6})$ and a good collapse is observed. Previous studies have found that the length scale of the magnetic field varies according to $Rm^{-1/2}$ (e.g. Christensen & Tilgner, 2004; Soderlund et al., 2015; Menu et al., 2020). We can relate this scaling to the Ekman number via
@@
since $Pm=O(1)$ and $Re=O(Ek^{-1/3})$ in our cases. We note that whereas the $O(Ek^{1/6})$ scaling in the magnetic Taylor microscale is weaker than the $O(Ek^{1/3})$ scaling in the velocity Taylor microscale, the computed values for both of these length scales are comparable over the parameter regime covered here (cf. Aubert et al., 2017; Davidson, 2013).
@@ 5 Conclusion
Table 1: Summary of asymptotic scalings for various quantities. Second column: predictions from QG plane layer theory (Calkins et al., 2015). Third column: observed non-magnetic scalings. Fourth column: observed dynamo scalings. These scalings assume the governing equations have been non-dimensionalised using the shell depth and the kinematic viscosity. The QG theory prediction for $B^{\prime}$ assumes PmPm is fixed.
@@
Understanding the dynamics of Earth’s outer core and the interiors of other planets and stars requires the extrapolation of model output by many orders of magnitude in parameter space. The identification of asymptotic behaviour in model output is paramount for this extrapolation process. A key component of this process is the identification of the asymptotically small (or large) parameter. In non-magnetic rotating convection this parameter is the Ekman number, $Ek$, and all other small parameters present, including the Rossby number, can be directly tied to $Ek$ via so-called distinguished limits. The dynamics of rapidly rotating convection is QG; a rigorous derivation of the nonlinear three-dimensional QG model is possible in planar (Sprague et al., 2006), and annular geometries with small radial length scales (Calkins et al., 2013). Recent work shows that the dominant scalings characterising the small-scale dynamics in both planar and spherical geometries are identical when the magnetic field is either not sufficiently strong to disrupt the geostrophic balance (Yan & Calkins, 2022a, b) or is not present at all (Nicoski et al., 2024). Given that the force balance present within the outer core is not directly observable, it is important to determine how the magnetic field influences the dynamics and resulting scaling behaviour as it becomes increasingly strong. Towards this end, we have used a set of numerical models that exhibit a leading-order MAC balance to determine how the dynamics depend on the Ekman number, and compared the results with equivalent simulations in which no magnetic field is present. Our findings, given in terms of the asymptotic scalings, are summarized in table 1. These results help to explain the behaviour of the forces studied in prior work (e.g. Soderlund et al., 2015; Yadav et al., 2016), and therefore provide a quantitative measure for the degree of asymptoticness that is obtained in any given simulation.
@@
The dynamo cases exhibit a MAC force balance for sufficiently large values of $Pm$, with the Coriolis force, pressure force, buoyancy force, and Lorentz force entering at leading order. The Lorentz force was observed to follow a stronger scaling with the Ekman number than the Coriolis force and buoyancy force. This finding suggests that either the Lorentz force becomes asymptotically larger than the Coriolis force as the Ekman number is decreased, in which case the Lorentz force would have to be balanced by the pressure gradient force, or that the scaling must change at smaller Ekman number beyond the current reach of simulations. The buoyancy force is found to be approximately the same order as the Coriolis force as $\widetilde{Ra}$ is increased, which, when combined with the approximately constant fluctuating temperature, leads to a simple relation for the Reynolds number. Notably, this scaling relationship, $Re_{c}\sim RaEk/Pr$, is equivalent to that derived from the CIA balance (e.g. Aurnou et al., 2020), despite the fact that the balances used to derive the relationship occur at different asymptotic orders. The order unity temperature fluctuation in the dynamos implies that the buoyancy force is the same asymptotic order as the Coriolis force whereas the buoyancy force for the non-magnetic cases is asymptotically smaller than the Coriolis force; this asymptotic difference in buoyancy force is the most notable difference between the dynamo and non-magnetic cases. We have argued that this difference in asymptotic order likely arises from the ohmic dissipation term in the dissipation equation.
@@
The magnetic Prandtl number plays a critical role in dynamos since, for a fixed value of buoyancy forcing ($\widetilde{Ra}$), it controls the relative size of the Lorentz force and therefore the observed force balance. Previous work has often used the strategy of employing as small a value of $Pm$ as can be used to generate dynamo action, such that the Lorentz force might play a secondary role. The majority of the simulations used in the present work have focused on a relatively large value of $Pm=2$, since if changes in the asymptotic behaviour occur due to the influence of magnetic fields, then such changes should be more readily observable with strong magnetic fields. For sufficiently small $Pm$, studies of plane layer dynamos find that the buoyancy force and Lorentz force enter at approximately the same order as the viscous force (Yan & Calkins, 2022a), whereas in the present study at constant $Pm$ both the buoyancy force and Lorentz force enter at an asymptotic order at least as large as the Coriolis force. It is possible that these two forces can enter at any asymptotic order between the viscous force and Coriolis force for an appropriate choice of $Pm$. The advection term, viscous force, and Coriolis force appear to follow the same asymptotic scaling at the fixed value of $Pm$ used in this paper as was found in the low $Pm$ values of Yan & Calkins (2022a), which suggests these terms do not depend asymptotically on $Pm$, although they are not independent of $Pm$.

### v3

@@ 4.1 Parameter regime and dependence on the magnetic Prandtl number
As the force balance is known to vary with $Pm$, we start by showing the radial fluctuating force balance as a function of $Pm$ for the case $Ek=10^{-5}$ and $Ra=1.5\times 10^{8}$ in figure 1(a). Note that when all other parameters are held fixed, the Lorentz force is controlled only by $Pm$. The force balance for a non-magnetic case is shown to the left of the dashed line. At small values of $Pm$, the force balance for the dynamo cases is similar to the non-magnetic case. As expected, the Lorentz force increases with $Pm$ [eD16, tS19, mM20]. At the largest value of $Pm=5$ reached here, the Lorentz force is slightly larger than the Coriolis force. The buoyancy force also increases with $Pm$, although less strongly than the Lorentz force – the reason for this increase is discussed later. The remaining terms tend to decrease as $Pm$ is increased, which is likely due to a decrease in flow speeds as the dynamo becomes more efficient at converting kinetic energy to magnetic energy. It is therefore apparent that the dynamo cases have force balances ranging from QG to MAC depending on the value of $Pm$, which suggests that $Pm$ can affect the asymptotic order that the various forces enter in the momentum equation. In this paper, we will fix $Pm=2$, which leads to cases with a MAC force balance, especially at small Ekman numbers. Thus, these results can be contrasted with prior work on the asymptotics of small magnetic Prandtl number dynamos, which are essentially QG [mY22].
@@ 4.2 Flow speeds
As indicated in figure 2(b), the dynamo cases are described well by the $O(Ek^{-1/3})$ scaling predicted from QG theory. The non-magnetic data appears to be less well described when viewed in these coordinates. To better understand this difference, we also plot the rescaled Reynolds number for the non-magnetic cases as a function of $Ra/Ra_{c}$ in figure 2(c), where good collapse is observed, indicating that the flow speeds are accurately described by the $O(Ek^{-1/3})$ scaling. Here, values of $Ra_{c}$ are taken from [uC06]. We believe that the better collapse in this view of parameter space is because $Ra_{c}$ captures additional effects that may be important for the non-magnetic data, as opposed to using $\widetilde{Ra}$ alone. In particular, as described in [eD04], the critical Rayleigh number can be represented in terms of a composite asymptotic expansion, consisting of the well-known $Ek^{-4/3}$ leading order term, along with corrections of order $O(Ek^{-4/3}Ek^{2/9})=O(Ek^{-10/9})$ and $O(Ek^{-4/3}Ek^{1/6})=O(Ek^{-7/6})$ (see their equations 3.25), that arise from the effects of the long radial modulation scale and Ekman pumping, respectively. Why the dynamos do not seem to be influenced by these corrections remains unclear, though the magnetic data exhibits clear evidence for asymptotic scaling, even at relatively modest Ekman numbers.
Noting the near linear trend of the convective Reynolds number for the dynamo cases in figure 2(b), we plot the convective Reynolds number compensated by $\widetilde{Ra}\vphantom{Ra}^{-1}$ in figure 2(d). The compensated data has much less dependence on $\widetilde{Ra}$, at least for sufficiently large values of $\widetilde{Ra}$. This scaling can be derived by balancing the Coriolis and buoyancy forces such that
@@
The reduced zonal Reynolds number is shown in figure 2(e), where again the dynamo data appears better collapsed in comparison to the non-magnetic data. If we again assume a balance between the Coriolis and buoyancy forces, but now in the mean momentum equation, then we obtain an identical scaling to that given above for the convective Reynolds number, i.e. $\overline{Re}_{\phi}\sim EkRa/Pr$. The compensated data given in figure 2(f) indicates that the dynamo cases are well described with this scaling, whereas the non-magnetic cases are not. The mean force balance in dynamos is known to be well described by the thermal wind balance when the Lorentz force does not enter the leading order force balance [jA05, mC21]. The mean force balance in the $\phi$-direction is between the Lorentz and Coriolis forces, though these components tend to be weaker in comparison to the leading order forces in the radial and colatitudinal directions [jA05, mC21, rO21, yL25]. However, for non-magnetic convection it is possible for the mean flows to be in geostrophic balance for sufficiently small Ekman number so that only large-scale viscous diffusion is available to saturate the zonal flow, as shown for non-magnetic convection using stress-free mechanical boundary conditions [jN24].
@@ 4.3 Force scalings
The rescaled Lorentz force is shown in figure 6(a), where an empirical scaling of $O(Ek^{-3/2})$ has been used to collapse the data. The ratio of the Lorentz force to the Coriolis force is shown in figure 6(b), where it can be seen that this ratio increases with decreasing $Ek$. However, note that while the Lorentz force is scaling more strongly with Ekman number than the Coriolis force, the Coriolis force is larger than the Lorentz force for most of the cases. It is important to recall that $Pm$ is constant in these cases; as shown in figure 1, the strength of the Lorentz force depends strongly on $Pm$. We also note that this scaling of the Lorentz force together with the scaling of the length scale of the Lorentz force of $O(Ek^{1/6})$ (shown later) implies that the magnetic field scales as $O(Ek^{-1/6})$, which is different than the order one scaling proposed in [mC21]. A plot of the rms of the fluctuating magnetic field is provided in figure 6(c), and the corresponding rescaled data is given in figure 6(d).
@@
The temporal behaviour of the fluctuating forces is shown in figure 7 for both a non-magnetic and dynamo case at $Ek=10^{-5}$, $Ra=1.5\times 10^{8}$ ($\widetilde{Ra}=32.3$) at a fixed point in the equatorial plane that lies midway between the shells. In our non-dimensionalisation, time is in units of the large-scale viscous diffusion time, $t_{\nu}$. However, we choose to plot the data in terms of the small-scale viscous diffusion time $Ek^{-2/3}t_{\nu}$, which is the dynamically relevant timescale for QG convection [tO23, tO25, e.g.]. For the non-magnetic case, the leading-order force balance is between the Coriolis force and the pressure gradient force, as would be expected for QG convection. The dynamo case is well described by a MAC balance in which the Coriolis force, pressure force, buoyancy force, and Lorentz force are all comparable and leading-order in the dynamics. Although these four terms tend to be the largest in a time averaged sense, the force balance exhibits significant changes with time. For example, at time $Ek^{-2/3}t_{\nu}\approx 6$, there is an approximate balance between the Lorentz and pressure forces, while at other times the Coriolis force is larger than the Lorentz force. Therefore, the leading-order force balance for the non-magnetic cases and the dynamo cases is different, as suggested by the asymptotic scalings for these forces. We note that other studies find the force balance might depend on length scale [tS19, jA17, tT21, e.g.], with many studies finding that the QG balance holds only at large length scales for dynamo simulations. However, the data shown here indicates that the pointwise balance is not geostrophic.
Further information on the spatial variation of these forces is provided in figure 8, which shows the radial fluctuating forces as a function of latitude for the same cases as in figure 7. For the non-magnetic case, a dominant balance between the Coriolis force and pressure gradient force exists across almost all latitudes. The advection term is larger than the viscous term outside the tangent cylinder, whereas the opposite is true inside the tangent cylinder. This difference is likely due to the higher Rayleigh numbers needed for convection within the tangent cylinder [tG23, e.g.]. For the dynamo case, a dominant balance between the Coriolis force, pressure gradient force, Lorentz force, and buoyancy force prevails across all latitudes, except near the poles where the Coriolis force becomes small. Therefore, for both the dynamo and non-magnetic cases, the dominant force balance does not vary strongly with latitude, although the absolute sizes of the various terms are sensitive to latitude. Most notably, there is a decrease in the size of the Coriolis, pressure gradient, viscous, and buoyancy forces by a few orders of magnitude inside the tangent cylinder for the non-magnetic case.
@@
As done in section 3, if we define $\tau$ as the relevant dynamical timescale and balance inertia with either of these terms then we have
@@
Interestingly, this is the asymptotic scaling for the period of convective Rossby waves, implying that these dynamics persist despite the leading order MAC balance.
@@ 4.4 Temperature and dissipation relations
Analysis of the forces in section 4.3 shows that the fluctuating buoyancy force, and therefore the fluctuating temperature, is characterised by a unique asymptotic scaling depending on whether the magnetic field is present. Figure 9 shows the rms of the fluctuating temperature for all cases, where this difference can be observed – for non-magnetic rotating convection the magnitude of the fluctuating temperature scales according to $Ek^{1/3}$, whereas the dynamo cases exhibit a fluctuating temperature that becomes independent of both $Ek$ and $\widetilde{Ra}$, as has been noted previously [mC21]. This difference can be explained based on energetic arguments as follows. For the dynamo cases, a dissipation equation can be derived by dotting the full momentum equation with the velocity field and averaging over both the fluid volume ($V$) and time to give
$\int_{V}$ $u_{r}f_{b}\,dV=\int_{V}u^{\prime}_{r}f^{\prime}_{b}+\overline{u}_{r}\overline{f}_{b}\,dV=\int_{V}\bm{\omega}^{2}\,dV$ (27)
@@
$\langle u^{\prime}_{r}f^{\prime}_{b}\rangle\sim\langle\bm{\omega}^{2}\rangle=\langle\mathbf{u}\cdot\mathbf{f}_{v}\rangle,$ (28)
@@ 4.5 Length scales
In this section we compare various length scales for both the dynamo and non-magnetic cases. Figure 10 shows visualisations of both the velocity field and the fluctuating temperature field. As noted in previous work [rY16, tS19, e.g.], the dynamo cases can form length scales that are larger than those observed in comparison to the non-magnetic cases, although small length scales are still present in the dynamo cases. For example, localised small-scale fluid structures with significant flow speeds can be seen in figure 10(b), especially near the inner and outer boundaries. This observation suggests that the dynamo simulations have a wider range of energetically relevant length scales than the non-magnetic cases. Note also the general lack of axial alignment in the dynamo cases, which likely results from the Lorentz force breaking the balance between the Coriolis force and the pressure gradient force.
@@
The linear asymptotic theory of rapidly rotating convection shows that, in addition to the $Ek^{1/3}$ ‘fast’ scale, a ‘slow’ radial scale is also present with an asymptotic scaling of $Ek^{2/9}$ and $Ek^{1/6}$ for the spherical shell and full sphere geometries, respectively [cJ00, eD04]. Thus, it might be anticipated that different regions of the spectra should be characterised by different asymptotic scalings. For our investigated range of parameters there is very little difference in these two scaling exponents so that $Ek^{1/6}$ was used here for simplicity. As shown later, this larger length scale may be pertinent to the generation of the magnetic field.
@@
A comparison of the velocity Taylor microscale for the dynamo and non-magnetic cases is shown in figure 13(c). The length scale for the non-magnetic cases tends to be slightly smaller than for the dynamo cases, as might be expected based on the visualisations in figure 10. However, the difference in the Taylor microscale between the non-magnetic and dynamo cases is small, which might be due to viscous dissipation in the dynamo simulations being concentrated in small, localized regions. In addition, the Taylor microscale scales approximately as $O(Ek^{1/3})$ as shown in figure 13(d), although there is still a weak trend in the rescaled data for both the dynamo and non-magnetic cases. This trend might suggest a slow rate of convergence to the predicted QG scaling. Other studies have also found that the velocity field retains an $O(Ek^{1/3})$ length scale [eK13b], though we stress here that this scaling does not necessarily indicate that the viscous force is dominant since, as we have shown, the same length scale can be deduced from the advection term.
@@
Finally, two magnetic length scales are shown in figure 14, where both the magnetic Taylor microscale ($\lambda^{\prime}_{b}$) and a length scale based on the Lorentz force ($\ell^{\prime}_{b}$) are shown. These length scales are defined in equations (11) and (13). Both length scales are comparable in magnitude over the investigated range of parameters and both follow similar trends with $\widetilde{Ra}$ and $Ek$. In particular, for smaller values of $\widetilde{Ra}$ these length scales show an increase up to $\widetilde{Ra}\sim 20-30$, and then a monotonic decrease is observed as $\widetilde{Ra}$ is increased. Figure 14(b) shows the magnetic length scales rescaled with $O(Ek^{1/6})$ and a good collapse is observed. Previous studies have found that the length scale of the magnetic field varies according to $Rm^{-1/2}$ [uC04, kS15, mM20, e.g.]. We can relate this scaling to the Ekman number via
@@
since $Pm=O(1)$ and $Re=O(Ek^{-1/3})$ in our cases. We note that whereas the $O(Ek^{1/6})$ scaling in the magnetic Taylor microscale is weaker than the $O(Ek^{1/3})$ scaling in the velocity Taylor microscale, the computed values for both of these length scales are comparable over the parameter regime covered here [jA17, pD13b, cf.].
@@ 5 Conclusion
Table 1: Summary of asymptotic scalings for various quantities. Second column: predictions from QG plane layer theory [mC15b]. Third column: observed non-magnetic scalings. Fourth column: observed dynamo scalings. These scalings assume the governing equations have been non-dimensionalised using the shell depth and the kinematic viscosity. The QG theory prediction for $B^{\prime}$ assumes PmPm is fixed.
@@
Understanding the dynamics of Earth’s outer core and the interiors of other planets and stars requires the extrapolation of model output by many orders of magnitude in parameter space. The identification of asymptotic behaviour in model output is paramount for this extrapolation process. A key component of this process is the identification of the asymptotically small (or large) parameter. In non-magnetic rotating convection this parameter is the Ekman number, $Ek$, and all other small parameters present, including the Rossby number, can be directly tied to $Ek$ via so-called distinguished limits. The dynamics of rapidly rotating convection is QG; a rigorous derivation of the nonlinear three-dimensional QG model is possible in planar [mS06], and annular geometries with small radial length scales [mC13]. Recent work shows that the dominant scalings characterising the small-scale dynamics in both planar and spherical geometries are identical when the magnetic field is either not sufficiently strong to disrupt the geostrophic balance [mY22, mY22b] or is not present at all [jN24]. Given that the force balance present within the outer core is not directly observable, it is important to determine how the magnetic field influences the dynamics and resulting scaling behaviour as it becomes increasingly strong. Towards this end, we have used a set of numerical models with $Pr=1$, $Pm=2$ (with a subset of simulations varying $Pm$ from $0.2$ to $5$), $Ek\in\{3\times 10^{-6},10^{-5},3\times 10^{-5},10^{-4}\}$ and $Ra$ up to $50$ times supercritical to determine how the dynamics depend on the Ekman number, and compared the results with equivalent simulations in which no magnetic field is present. Our findings, given in terms of the asymptotic scalings, are summarized in table 1. These results help to explain the behaviour of the forces studied in prior work [kS15, rY16, e.g.], and therefore provide a quantitative measure for the degree of asymptoticness that is obtained in any given simulation.
@@
The dynamo cases exhibit a MAC force balance for sufficiently large values of $Pm$, with the Coriolis force, pressure force, buoyancy force, and Lorentz force entering at leading order. This large-scale balance has been observed in other direct numerical simulations that managed to perform a limited number of simulations at lower $Ek$ and $Pm$ than our model suite [nS17, yL25]. In our simulations, the Lorentz force was observed to follow a stronger scaling with the Ekman number than the Coriolis force and buoyancy force. This finding suggests that either the Lorentz force becomes asymptotically larger than the Coriolis force as the Ekman number is decreased, in which case the Lorentz force would have to be balanced by the pressure gradient force, or that the scaling must change at smaller Ekman number beyond the current reach of simulations. The buoyancy force is found to be approximately the same order as the Coriolis force as $\widetilde{Ra}$ is increased, which, when combined with the approximately constant fluctuating temperature, leads to a simple relation for the Reynolds number. Notably, this scaling relationship, $Re_{c}\sim RaEk/Pr$, is equivalent to that derived from the CIA balance [jmA20, e.g.], despite the fact that the balances used to derive the relationship occur at different asymptotic orders. The order unity temperature fluctuation in the dynamos implies that the buoyancy force is the same asymptotic order as the Coriolis force whereas the buoyancy force for the non-magnetic cases is asymptotically smaller than the Coriolis force; this asymptotic difference in buoyancy force is the most notable difference between the dynamo and non-magnetic cases. We have argued that this difference in asymptotic order likely arises from the ohmic dissipation term in the dissipation equation.
@@
The magnetic Prandtl number plays a critical role in dynamos since, for fixed values of $\widetilde{Ra}$ and $Ek$, it can be used to control the relative size of the Lorentz force and therefore the observed force balance. Previous work has often used the strategy of employing as small a value of $Pm$ as can be used to generate dynamo action, such that the Lorentz force might play a secondary role. The majority of the simulations used in the present work have focused on a relatively large value of $Pm=2$, since if changes in the asymptotic behaviour occur due to the influence of magnetic fields, then such changes should be more readily observable with strong magnetic fields. For sufficiently small $Pm$, studies of plane layer dynamos find that the buoyancy force and Lorentz force enter at approximately the same order as the viscous force [mY22], whereas in the present study at constant $Pm$ both the buoyancy force and Lorentz force enter at an asymptotic order at least as large as the Coriolis force. It is possible that these two forces can enter at any asymptotic order between the viscous force and Coriolis force for an appropriate choice of $Pm$. The advection term, viscous force, and Coriolis force appear to follow the same asymptotic scaling at the fixed value of $Pm$ used in this paper as was found in the low $Pm$ values of [mY22], which suggests these terms do not depend asymptotically on $Pm$, although they are not independent of $Pm$. Future work is needed to understand the dynamics in the dual limit $(Ek,Pm)\rightarrow 0$.